# MoEMba: A Mamba-based Mixture of Experts for High-Density EMG-based Hand Gesture Recognition

### v1

@@ I Introduction
In greater detail, a key challenge in practical implementations is the low accuracy of inter-session and inter-subject classification, with session-to-session variability reaching as high as 1010101040404040%, which arises from the inherent temporal variability of sEMG signals. This variability is influenced by several factors, including muscle condition (fatigue, atrophy, or hypertrophy), changes in skin impedance (such as sweating), anatomical differences among subjects, non-stationarities due to limb positioning and force variations, and electrode drift, all of which contribute to accuracy degradation. Moreover, these factors can introduce considerable inconsistencies in signal patterns across different sessions and individuals, thus complicating the generalization of gesture recognition algorithms [20, 30, 31, 32].
@@ A. Dataset
In this study, we used the CapgMyo HD-sEMG dataset [30], recorded from $128$ channels sampled at $1000$Hz. We utilized sub-databases DB-a ($23$ subjects performing $8$ gestures held for $3-10$s) and DB-b (a subset of DB-a comprising $10$ subjects with recordings from two sessions separated by a minimum of $7$ days, performing the same $8$ gestures with hold durations of $\sim$3333s). The CapgMyo HD-sEMG dataset configuration introduces inter-session and inter-subject variability, which challenge model robustness—an issue that this study aims to address. To capture richer semantic information and local features, the data was pre-processed using $45-55$Hz Butterworth filter, followed by segmentation into overlapping $64$ms windows with $8$ms steps, resulting in $64\times 128$ matrices, and finally normalized to $[-1,1]$.
@@ III The MoEMba Framework
Figure 1: MoEMba Framework. (a) Raw time series data is pre-processed via patching. (b) Shallow features are extracted using small $(3\times 3)$ and large $(7\times 7)$ receptive field convolutions, and wavelet transform-modulated features, with channel attention applied to EMG signal patches. (c) A Mixture of Experts (MoE) block routes patches to two Mamba experts based on a gating network. (d) The Mamba block incorporates projections, $1$D convolution, selective State-Space Modeling (SSM), and skip connections for temporal dependency learning.
@@ A. The WTFM Block
The MoEMba processes each signal patch, $P_{i}\in\mathbb{R}^{L\times V}$, by the WTFM block, where $L$ defines the temporal length and $V$ denotes the number of channels. This $2$D representation facilitates convolutional operations, allowing a more effective capture of spatial relationships across channels. The WTFM block initially extracts local and global temporal features using small (e.g., $3\times 3$) and large (e.g., $7\times 7$) temporal convolutional kernels, respectively. A Discrete Wavelet Transform (DWT) is then applied to the local temporal features (extracted by the small temporal convolutional kernel), decomposing them into four directional components: approximation $(c_{A})$, horizontal detail $(c_{H})$, vertical detail $(c_{V})$, and diagonal detail $(c_{D})$. These components are calculated based on different frequency bands of the wavelet decomposition. These components capture detailed temporal variations at different scales. DWT components are then upsampled back to the original size of $P_{i}$ using bicubic interpolation. These upsampled wavelet features are then refined using a channel attention mechanism, which learns to focus on the most important channels. The attention weights $\alpha_{i}$ are given by
@@ B. The MoE Block
We incorporate an adapted gating mechanism inspired by [28] that introduces optimizing enhanced sparsity to optimize computational efficiency of the underlying model. More specifically, a tunable sparsity mechanism is employed where Gaussian noise, controlled by a trainable weight matrix $W_{\text{noise}}$, is added to the network output before applying the Softmax function. To further promote sparsity, only the top $k$ values from the network output are retained, while the remaining values are suppressed to $\infty$- ∞, effectively setting their corresponding gate values to zero. This selective approach is formalized as follows
[…]
and $\text{TopKGating}(v,k)_{i}$ retains only the top $k$ elements of $v$, setting the rest to $\infty$- ∞. Here, $W_{\text{gate}}$ and $W_{\text{noise}}$ are trainable weight matrices. To prevent the potential bias in the gating network towards a single expert, a balanced gate mechanism is implemented. This is achieved by minimizing a balance loss $\mathcal{L}_{B}$, which encourages a more uniform distribution of samples across experts. The balance loss is defined as
@@ IV Experimental Results
All experiments were conducted on a computing system equipped with an NVIDIA GeForce RTX $2080$ Ti GPU, an Intel i9999-$9820$X CPU, and $128$GB of RAM. The MoEMba framework was configured with several key hyperparameters. The SSM employed a state expansion factor of $16$ and a local convolution width of $4$, with a model dimension of $128$. The MoE architecture consisted of two experts and two gating networks, with a hidden dimension factor of $1$. Each Mamba expert within the MoE framework used a block expansion factor of $4$. A small number of experts was chosen to reduce the risk of overfitting, particularly given the dataset size, while ensuring that each expert receives sufficient data for effective training. The model was trained over $50$ epochs, using a batch size of $32$ and an initial learning rate of $0.0001$, which was adjusted via a cosine annealing schedule.
@@
TD K. Englehart et al. [42] $0.416$±$0.198$
ETD R. N. Khushaba et al. [43] $0.437$±$0.209$
NinaPro M. Atzori et al. [44] $0.425$±$0.170$
SampEn A. Phinyomark et al. [45] $0.448$±$0.217$
TVGGNet J. Pereira et al. [29] $0.420$±$0.183$
The Proposed MoEMba This Work $0.569$±$0.201$

### v2

@@ I Introduction
In greater detail, a key challenge in practical implementations is the low accuracy of inter-session and inter-subject classification, with session-to-session variability reaching as high as 10104040%, which arises from the inherent temporal variability of sEMG signals. This variability is influenced by several factors, including muscle condition (fatigue, atrophy, or hypertrophy), changes in skin impedance (such as sweating), anatomical differences among subjects, non-stationarities due to limb positioning and force variations, and electrode drift, all of which contribute to accuracy degradation. Moreover, these factors can introduce considerable inconsistencies in signal patterns across different sessions and individuals, thus complicating the generalization of gesture recognition algorithms [26, 8, 25, 20].
@@ A. Dataset
In this study, we used the CapgMyo HD-sEMG dataset [8], recorded from $128$ channels sampled at 10001000Hz. We utilized sub-databases DB-a ($23$ subjects performing $8$ gestures held for 3103-10s) and DB-b (a subset of DB-a comprising $10$ subjects with recordings from two sessions separated by a minimum of $7$ days, performing the same $8$ gestures with hold durations of $\sim$33s). The CapgMyo HD-sEMG dataset configuration introduces inter-session and inter-subject variability, which challenge model robustness—an issue that this study aims to address. To capture richer semantic information and local features, the data was pre-processed using 455545-55Hz Butterworth filter, followed by segmentation into overlapping 6464ms windows with 88ms steps, resulting in $64\times 128$ matrices, and finally normalized to [1,1][-1,1].
@@ III The MoEMba Framework
Figure 1: MoEMba Framework. (a) Raw time series data is pre-processed via patching. (b) Shallow features are extracted using small $(3\times 3)$ and large $(7\times 7)$ receptive field convolutions, and wavelet transform-modulated features, with channel attention applied to EMG signal patches. (c) A Mixture of Experts (MoE) block routes patches to two Mamba experts based on a gating network. (d) The Mamba block incorporates projections, 11D convolution, selective State-Space Modeling (SSM), and skip connections for temporal dependency learning.
@@ A. The WTFM Block
The MoEMba processes each signal patch, $P_{i}\in\mathbb{R}^{L\times V}$, by the WTFM block, where $L$ defines the temporal length and $V$ denotes the number of channels. This 22D representation facilitates convolutional operations, allowing a more effective capture of spatial relationships across channels. The WTFM block initially extracts local and global temporal features using small (e.g., $3\times 3$) and large (e.g., $7\times 7$) temporal convolutional kernels, respectively. A Discrete Wavelet Transform (DWT) is then applied to the local temporal features (extracted by the small temporal convolutional kernel), decomposing them into four directional components: approximation $(c_{A})$, horizontal detail $(c_{H})$, vertical detail $(c_{V})$, and diagonal detail $(c_{D})$. These components are calculated based on different frequency bands of the wavelet decomposition. These components capture detailed temporal variations at different scales. DWT components are then upsampled back to the original size of $P_{i}$ using bicubic interpolation. These upsampled wavelet features are then refined using a channel attention mechanism, which learns to focus on the most important channels. The attention weights $\alpha_{i}$ are given by
@@ B. The MoE Block
We incorporate an adapted gating mechanism inspired by [3] that introduces optimizing enhanced sparsity to optimize computational efficiency of the underlying model. More specifically, a tunable sparsity mechanism is employed where Gaussian noise, controlled by a trainable weight matrix $W_{\text{noise}}$, is added to the network output before applying the Softmax function. To further promote sparsity, only the top $k$ values from the network output are retained, while the remaining values are suppressed to -$\infty$, effectively setting their corresponding gate values to zero. This selective approach is formalized as follows
$\!\!\!\!\!\!\!\!\mathcal{G}(Z_{p})$ $\!\!\!\!=\!\!\!\!$ $\text{Softmax}(\text{TopKGating}(H(Z_{p}),k)),$ (11)
$\!\!\!\!\!\!\!\!\text{where }H_{i}(Z_{p})$ $\!\!\!\!=\!\!\!\!$ $(Z_{p}W_{\text{gate}})_{i}+\text{Softplus}((Z_{p}W_{\text{noise}})_{i}),$ (12)
and $\text{TopKGating}(v,k)_{i}$ retains only the top $k$ elements of $v$, setting the rest to -$\infty$. Here, $W_{\text{gate}}$ and $W_{\text{noise}}$ are trainable weight matrices. To prevent the potential bias in the gating network towards a single expert, a balanced gate mechanism is implemented. This is achieved by minimizing a balance loss $\mathcal{L}_{B}$, which encourages a more uniform distribution of samples across experts. The balance loss is defined as
@@ IV Experimental Results
All experiments were conducted on a computing system equipped with an NVIDIA GeForce RTX $2080$ Ti GPU, an Intel i99-98209820X CPU, and 128128GB of RAM. The MoEMba framework was configured with several key hyperparameters. The SSM employed a state expansion factor of $16$ and a local convolution width of $4$, with a model dimension of $128$. The MoE architecture consisted of two experts and two gating networks, with a hidden dimension factor of $1$. Each Mamba expert within the MoE framework used a block expansion factor of $4$. A small number of experts was chosen to reduce the risk of overfitting, particularly given the dataset size, while ensuring that each expert receives sufficient data for effective training. The model was trained over $50$ epochs, using a batch size of $32$ and an initial learning rate of $0.0001$, which was adjusted via a cosine annealing schedule.
@@
TD K. Englehart et al. [10] 0.4160.416±$0.198$
ETD R. N. Khushaba et al. [24] 0.4370.437±$0.209$
NinaPro M. Atzori et al. [4] 0.4250.425±$0.170$
SampEn A. Phinyomark et al. [36] 0.4480.448±$0.217$
TVGGNet J. Pereira et al. [35] 0.4200.420±$0.183$
The Proposed MoEMba This Work 0.5690.569±$0.201$